Colin is the managerial accountant in charge of Company A, which
sells water bottles. He previously determined that the fixed costs of
Company A consist of property taxes, a lease, and executive salaries,
which add up to $100,000 per year. The variable cost associated with
producing one water bottle is $2 per unit. The water bottle is sold at a
premium price of $12.
a) Prepare B.E. chart. b) What volume per month is required in order to break even?
c) What profit would be realized on a monthly volume of 60,000
units?
d) What volume would be needed to obtain a profit of $ 20,000
per month?
e) What volume is required to provide revenue of $ 40,000 per
month?
1
Expert's answer
2020-04-08T09:48:49-0400

a) A break even chart is a chart that shows the sales volume level at which total costs equal sales.

b) The volume per month that is required in order to break even is:

"Q = \frac{FC} {P - AVC} = \frac{100,000}{12-2} = 10,000" units.

c) If Q = 60,000 units, then total profit is: "TP = P\u00d7Q - (FC + VC) = 12\u00d760,000 - (100,000 + 2\u00d760,000) = 500,000."

d) To obtain a profit of $ 20,000 per month the quantity is:

"Q = \frac{TP + FC} {P - AVC} = \frac{20,000 + 100,000} {12 - 2} = 12,000"units12,000units.

e) The volume required to provide revenue of $ 40,000 per month is:

"Q = TR\/P = 40,000\/12 = 3,333 units."